OpenStudy (anonymous):

Solve the linear system. x + y = -3 y = 2x

OpenStudy (anonymous):

x + y = -3 y = 2x x + 2x = -3 3x = -3 x = -1

OpenStudy (anonymous):

And thus y = -2.

OpenStudy (anonymous):

\[\large x + y = -3\]\[\large y = 2x\] since \(y=2x\) we can substitute it into the equation \(x+y=-3\):\[x+(2x)=-3\]\[x+2x=-3\]combine like terms:\[3x=-3\]divide oth sides by \(3\) to isolate \(x\):\[x=-1\] now that we have a value for \(x\), we can plug it into the equation \(y=2x\) to solve for \(y\):\[y=2(-1)\]\[y=-2\] thus the solution to this system of equations is \((-1,-2)\)